Reduce the quadratic form 2𝑥

2 + 2𝑦

2 + 2𝑧

2 + 2𝑦𝑧 to the canonical form by 

orthogonal reduction. Find the index, signature and nature of the quadratic form.


1
Expert's answer
2022-02-08T12:30:59-0500

Matrix A"=\begin{pmatrix}
 2&0&0 \\
 0&2&1 \\
0&1&2
\end{pmatrix}"



"\>\>\begin{vmatrix}
 A-\lambda\>I \\
 
\end{vmatrix}=\begin{vmatrix}
 2-\lambda&0&0 \\
 0&2-\lambda& 1\\
0&1&2-\lambda
\end{vmatrix}=0"



"(2-\lambda)\begin{bmatrix}
 (2-\lambda)(2-\lambda)-1\\
 
\end{bmatrix}+0+0=0"


"(2-\lambda)(\lambda^2-4\lambda+3)=0"

"(2-\lambda)(\lambda-1)(\lambda-3)=0"

"\lambda=1,\lambda=2,\lambda=3"


For "\lambda=1,\begin{pmatrix}
 1&0&0 \\
 0&1&1 \\
0&1&1
\end{pmatrix}\begin{pmatrix}
 x_1 \\
 x_2 \\
x_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0\\
0
\end{pmatrix}"


"x_1=0,\>x_2=-1,\>x_3=1"



For "\lambda=2\begin{pmatrix}
 0&0 & 0 \\
 0&0 & 1\\
0&1&0
\end{pmatrix}\begin{pmatrix}
 x_1 \\
 x_2 \\
x_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\
0
\end{pmatrix}"


"x_1=1,\>x_2=0,\>x_3=0"


For "\lambda=3\begin{pmatrix}
 -1&0&0 \\
 0&-1 & 1\\
0&1&-1
\end{pmatrix}\begin{pmatrix}
 x_1 \\
 x_2\\
x_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0\\
0
\end{pmatrix}"


"x_1=0,\>x_2=1,\>x_3=1"



Modal matrix "\begin{pmatrix}
 0&1&0 \\
 -1&0&1\\
1&0&1
\end{pmatrix}"


Normalized modal matrix

"Q=\begin{pmatrix}
 0&1&0 \\
 \frac{-1}{\sqrt2} & 0&\frac{1}{\sqrt2}\\
\frac{1}{\sqrt2}&0&\frac{1}{\sqrt2}
\end{pmatrix}"



Diagonalizing matrix

"D=Q^TAQ=\begin{pmatrix}
 0&\frac{-1}{\sqrt2}& \frac{1}{\sqrt2} \\
 1&0&0 \\
0&\frac{1}{\sqrt2}&\frac{1}{\sqrt2}
\end{pmatrix}\begin{pmatrix}
 2&0&0 \\
 0&2 & 1\\
0&1&2
\end{pmatrix}\begin{pmatrix}
 0&1&0 \\
 \frac{-1}{\sqrt2}&0&\frac{1}{\sqrt2} \\
\frac{1}{\sqrt2}&0&\frac{1}{\sqrt2}
\end{pmatrix}"



"=\begin{pmatrix}
 1&0&0 \\
 0&2& 0\\
0&0&3
\end{pmatrix}"


Diagonalized matrix has principal diagonal element =Eigenvalues

All other elements"=0"


The orthogonal transformation reduces the quadratic form to conical form

"y_1^2+2y_2^2+3y_3^2"


Index"=3"

Signature"=2\u00d73-3=3"

Nature of quadratic form

Positive definate